OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find 2·f(4).

OpenStudy (anonymous):

f(4)=(4^2)+1=9 2.f(4)= 2x9=18

OpenStudy (anonymous):

f(4)=(4^2)+1=16+1=17 2.f(4)=2*17=34

OpenStudy (anonymous):

ouch careless mistake :S